OpenStudy (anonymous):

A wheel 25.0cm in radius turning at 120rpm increases its frequency to 660rpm in 9.00s. Find (a) the constant angular acceleration in rad/s^2; and (b) the tangential acceleration of a point on it's rim.

OpenStudy (anonymous):

\[\alpha={{\Delta \omega} \over{\Delta t}}\] = (660 - 120)/9 = ? rad/s^2 tangential acceleration = acceleration x distance from center = the value above * 0.25 m = ?